OpenStudy (anonymous):

Find the inverse of the function. f(x) = x3 + 8 would this be y= 3^ √(x-8) or not a one to one function???

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

so y= 3^ √(x-8) is correct?

OpenStudy (jdoe0001):

\( y = x^3+8\\ \text{inverse will be, by swapping the variables}\\ x = y^3+8 \implies x-8 = y^3 \implies \huge \sqrt[3]{x-8} = y\)

OpenStudy (anonymous):

thankyou

OpenStudy (jdoe0001):

yw